# Cross-modal Full-mode Fine-grained Alignment for Text-to-Image Person Retrieval

### v1

@@ 4.2. Comparison with State-of-the-Art Methods
P+G UniPT (Shao et al., 2023) + IRRA (Jiang and Ye, 2023) 64.50 80.24 85.74 38.22
PLIP (Zuo et al., 2024) + IRRA (Jiang and Ye, 2023) 65.79 81.94 87.32 39.43
NAM (Tan et al., 2024) + IRRAR 66.34 81.94 86.73 40.14
NAM (Tan et al., 2024) + FMFA (ours) 66.58 81.94 87.04 40.17
HAM (Jiang et al., 2025) + IRRAR 68.21 83.28 88.04 41.72
HAM (Jiang et al., 2025) + FMFA (ours) 68.37 83.28 88.10 41.76
Performance Comparisons on ICFG-PEDES. We also evaluate FMFA on another benchmark, ICFG-PEDES, as shown in Table 4. Using VL-Backbones without ReID-domain pre-training, FMFA achieves the best performance across all evaluation metrics, attaining 64.29% Rank-1 and 39.43% mAP. Compared with IRRA, FMFA shows a notable improvement of 0.81% in Rank-1 and 1.23% in mAP, which is meaningful for practical applications. When adopting VL-Backbones with ReID-domain pre-training, FMFA yields slight gains, outperforming IRRA by 0.24% and 0.16% in Rank-1 with the NAM-based and HAM-based backbones, respectively.
In summary, our FMFA achieves the best performance across all evaluation metrics on three common public benchmarks. To the best of our knowledge, FMFA is the best method for all global matching methods. This demonstrates the generalizability and robustness of our proposed method.
Table 5. Ablation study on each component of FMFA on CUHK-PEDES, ICFG-PEDES and RSTPReid.
@@
SDM A-SDM EFA Rank-1 Rank-5 Rank-10 Rank-1 Rank-5 Rank-10 Rank-1 Rank-5 Rank-10
0 Baseline $\checkmark$ 73.45 89.38 93.69 63.48 80.16 85.78 59.50 81.80 88.85
1 +A-SDM $\checkmark$ 74.04 89.86 93.89 64.26 80.59 85.90 60.25 81.45 88.70
2 +EFA $\checkmark$ $\checkmark$ 73.73 89.64 94.04 63.77 80.39 85.86 60.45 82.30 89.25
3 FMFA $\checkmark$ $\checkmark$ 74.16 90.12 94.10 64.29 80.48 85.93 61.05 83.85 89.80
4.3. Ablation Study
In this section, we analyze our proposed components in the FMFA framework. For simplicity, we omit the components of $\mathcal{L}_{id}$ and the IRR module that were proposed by IRRA and used in all experiments. Only one of SDM and A-SDM can be used at the same time.
To fully demonstrate the impact of our proposed components in FMFA, we conduct an empirical analysis on three common datasets (i.e., CHUK-PEDES (Li et al., 2017), ICFG-PEDES (Ding et al., 2021) and RSTPReid (Zhu et al., 2021)). The Rank-1, Rank-5 and Rank-10 accuracies(%) are shown in Table 5.
Effect of The A-SDM Module. To evaluate the contribution of the Adaptive Similarity Distribution Matching (A-SDM) module, we perform ablation experiments by replacing the A-SDM module with the SDM module, keeping all hyperparameters unchanged. Specifically, as shown in Table 5, replacing A-SDM with SDM leads to a decrease in the Rank-1 accuracy by 0.59%, 0.78%, and 0.75% across the three datasets, as observed in No. 0 vs. No. 1. Additionally, all evaluation metrics on the CUHK-PEDES and ICFG-PEDES datasets degrade, further confirming the superiority of A-SDM. Moreover, when combined with the EFA module, the advantage of A-SDM becomes even more pronounced. As shown in No. 2 vs. No. 3, replacing the A-SDM module with the SDM module results in 1.26% and 1.00% decrease in Rank-1 and Rank-5 on the CUHK-PEDES dataset, respectively, and a 1.55% drop in the Rank-5 accuracy on the RSTPReid dataset. These results collectively validate the consistent and significant contribution of the A-SDM module to overall performance.
Effect of The EFA Module. To enhance the global matching capability of model, the Explicit Fine-grained Alignment (EFA) module introduces fine-grained cross-modal interaction based on a sparse similarity matrix. The effectiveness of the EFA module is demonstrated through the experimental comparisons between No. 0 vs. No. 2 and No. 1 vs. No. 3. Specifically, as shown in No. 1 vs. No. 3, removing the EFA module from FMFA leads to a performance drop of 0.26% and 0.21% in Rank-5 and Rank-10 on the CUHK-PEDES dataset, and a more significant decline of 2.40% and 1.10% in Rank-5 and Rank-10 on the RSTPReid dataset. However, EFA causes a 0.11% drop in Rank-5 on ICFG-PEDES, suggesting that its sparse and hard coding strategy, which focuses only on the most relevant patches, may overlook other informative ones and lead to information loss. Notably, this comparison also reflects the joint ablation of EFA and the A-SDM module, further verifying the complementary effect between the two modules. Moreover, to further validate the individual contribution of the EFA module, we design an additional experiment that retains the SDM module while removing only the EFA module, as observed in No. 0 vs. No. 2. In this setting, the absence of EFA results in 0.38%, 0.29%, and 0.95% drops in Rank-1 on CUHK-PEDES, ICFG-PEDES and RSTPReid, respectively. These comparisons demonstrate the effectiveness of the EFA module.
@@ 4.3. Ablation Study
4.4. Parameter Study
We conduct the parameter analysis on the CUHK-PEDES dataset, including three hyperparameters (i.e., the weight factor $\alpha$, the sparsity threshold $\sigma$ and the factor $\lambda$) and the weight of the proposed loss functions (i.e., $\mathcal{L}_{A-sdm}$ and $\mathcal{L}_{efa}$). While we analyze the special parameter, other parameters are kept the same as those described in Section 4.1.
Hyperparameters Analysis. As shown in Figure 4(a), we vary the weight factor $\alpha$ from 1 to 20. The results on the CUHK-PEDES dataset indicate that setting $\alpha$ to 10 yields the best performance across all metrics, suggesting that the adaptive pull force on unmatched positive pairs is optimal. However, when $\alpha$ is increased to 20, the performance drops significantly because the pull force becomes excessively strong, causing unmatched positive pairs to fail to align properly and instead over-pulling mismatched positives, which leads to false positives. We further vary the sparsity threshold $\sigma$ from $0.1/N$ to $20/N$ and the LSE factor $\lambda$ from 0.1 to 10, where $N$ denotes the number of image patches. As illustrated in Figure 4(b) and Figure 4(c), setting $\sigma$ to $1/N$ and $\lambda$ to 1 yields the best overall performance.
Loss Function Weights. We conduct a series of experiments on the CUHK-PEDES dataset to analyze the effect of the weights of the two proposed loss functions, $\mathcal{L}_{A-sdm}$ and $\mathcal{L}_{efa}$, varying them from 0.1 to 10.0, as shown in Table 6. The results show that setting both weights to 1 yields the best performance. Notably, increasing the weight of $\mathcal{L}_{efa}$ to 5.0 or 10.0 results in gradient explosion, making the model untrainable and causing all metrics to drop below 1.
Table 6. The sensitivity analysis of the weights of $\mathcal{L}_{A-sdm}$ and $\mathcal{L}_{efa}$ on the CUHK-PEDES dataset.
@@ 4.5. Qualitative Results
Visualization of The Sparse Process. We visualize the similarity maps before and after the sparse process (i.e., Eq. (5) and Eq. (6)) in the EFA module, as shown in Figure 5. Patches with high similarity (greater than 0.75) are preserved, whereas those with low similarity (below 0.5) are suppressed to values under 0.25 and thus omitted during aggregation (left in Figure 5). When most patches in an image exhibit relatively high similarity (around 0.75), the sparse process retains only the patches with the highest similarity, while pushing the similarity of the remaining patches below 0.5 (right in Figure 5). This ensures that the subsequent aggregation focuses only on the most relevant patches, reducing computational and memory costs while maintaining performance. Unlike implicit aggregation methods based on attention mechanisms, EFA explicitly aggregates image patches and text tokens, allowing us to observe whether the most relevant patches are effectively aggregated by visualizing the similarity between patches and tokens.
Inference Time Comparison We compare the inference time between FMFA and recent local matching methods (e.g., PLOT (Park et al., 2024) and RaSa (Bai et al., 2023b)) on the test sets of three datasets, as shown in Table 7. As a global matching method, FMFA only computes global features during inference, thus achieving a higher inference speed than local matching methods. Moreover, the efficiency advantage of FMFA becomes more pronounced as the size of the test set increases. For instance, the inference time of FMFA vs. PLOT increases from 3s vs. 5s on RSTPReid to 50s vs. 91s on ICFG-PEDES.
Table 7. Comparison of the inference time (s) between FMFA and recent local matching methods.
@@
PLOT (Park et al., 2024) 5 16 91
RaSa (Bai et al., 2023b) 388 1168 3871
@@
Figure 6. Comparison of top-5 retrieved results on CUHK-PEDES between baseline and FMFA for each text query. The image corresponding to query text, along with the matched and mismatched images, are highlighted with black, green and red rectangles, respectively.
Visualization of Top-5 Retrieval Results. Figure 6 compares the top-5 retrieval results between the baseline IRRAR and our proposed FMFA on the CUHK-PEDES dataset. As shown in Figure 6, FMFA achieves more accurate retrieval results, successfully retrieving images that the baseline fails to match. For query texts where the baseline performs well, FMFA further improves performance by retrieving more relevant pedestrian images (e.g., Figure 6(a) and Figure 6(c)). Even for hard negative samples, where the baseline struggles to retrieve the correct image, FMFA still enhances the similarity between positive pairs (e.g., Figure 6(b) and Figure 6(d)). This is because our proposed FMFA focuses on the unmatched positive pairs and adaptively pulls the positive pairs closer. More comparisons of the top-5 retrieved results between the baseline and FMFA on the ICFG-PEDES and RSTPReid datasets are provided in Figure 7 and Figure 8.
@@
Figure 7. Comparison of top-5 retrieved results on ICFG-PEDES between baseline and FMFA for each text query. The image corresponding to query text, matched and mismatched images are marked with black, green and red rectangles, respectively.
@@
Figure 8. Comparison of top-5 retrieved results on RSTPReid between baseline and FMFA for each text query. The image corresponding to query text, matched and mismatched images are marked with black, green and red rectangles, respectively.
5. Conclusion
In this paper, we propose a cross-modal Full-Mode Fine-grained Alignment (FMFA) framework to learn discriminative global text-image representations through full-mode fine-grained alignment, including explicit fine-grained alignment and existing implicit relational reasoning. We design an Adaptive Similarity Distribution Matching (A-SDM) module to concentrate on unmatched positive pairs, adaptively pulling them closer. In addition, to achieve cross-modal fine-grained alignment, we introduce an Explicit Fine-grained Alignment (EFA) module, which explicitly aggregates local text and image representations based on the sparse similarity matrix and employs a hard coding method. These modules work collaboratively to align images and text into a joint embedding space. Extensive experiments on three datasets demonstrate the superiority and effectiveness of our proposed FMFA framework.
Limitation. The fixed threshold in the sparse process only keeps the most relevant patches, which may result in the loss of semantic information and limit the effective aggregation of local features, thereby affecting the overall performance of our model. Incorporating adaptive methods that capture complete semantic information (e.g., tree transformer (Wang et al., 2019)) could further enhance our model.
Acknowledgements.
This work was supported by the National Natural Science Foundation of China (No. 62302080), National Foreign Expert Project of China (No. S20240327), Sichuan Science and Technology Program (No. 2025HJRC0021) and Sichuan Province Innovative Talent Funding Project for Postdoctoral Fellows (No. BX202312).

### v2

@@ 4.2. Comparison with State-of-the-Art Methods
P+G UniPT (shao2023unified) + IRRA (jiang2023cross) 64.50 80.24 85.74 38.22
PLIP (zuo2024plip) + IRRA (jiang2023cross) 65.79 81.94 87.32 39.43
NAM (tan2024harnessing) + IRRAR 66.34 81.94 86.73 40.14
NAM (tan2024harnessing) + FMFA (ours) 66.58 81.94 87.04 40.17
HAM (jiang2025modeling) + IRRAR 68.21 83.28 88.04 41.72
HAM (jiang2025modeling) + FMFA (ours) 68.37 83.28 88.10 41.76
Evaluation Results on ICFG-PEDES. We assess FMFA on the ICFG-PEDES benchmark, with the results displayed in Table 5. Using VL-Backbones without ReID-domain pre-training, FMFA obtains the leading results across all metrics, attaining 64.29% Rank-1 and 39.43% mAP. Compared with IRRA, FMFA shows a notable improvement of 0.81% Rank-1 and 1.23% mAP, which is meaningful for practical applications. When adopting VL-Backbones with ReID-domain pre-training, FMFA yields slight gains, outperforming IRRA by 0.24% and 0.16% in Rank-1 with the NAM-based and HAM-based backbones, respectively.
In conclusion, FMFA attains the highest performance across all evaluation metrics on the three widely used public benchmarks. As far as we are aware, FMFA is the best method for all global matching methods. This highlights the ability of our method to generalize well and maintain robustness.
4.3. Ablation Study
In this subsection, we examine our proposed components in the FMFA framework. For simplicity, we omit the components of $\mathcal{L}_{id}$ and the IRR module that were proposed by IRRA and used in all experiments. Only one of SDM and A-SDM can be used at the same time.
To thoroughly assess the contribution of our FMFA modules, we undertake an empirical analysis on three widely used datasets. Table 6 summarizes the Rank-1/5/10 accuracies (%) together with the mAP (%) performance.
Effect of The A-SDM Module. To evaluate the contribution of the Adaptive Similarity Distribution Matching (A-SDM) module, we perform ablation experiments by replacing the A-SDM module with the SDM module, keeping all hyperparameters unchanged. Specifically, as shown in Table 6, replacing A-SDM with SDM results in a reduction of Rank-1 accuracy by 0.59%, 0.78%, and 0.75% across the three datasets, and also causes a 1.19% drop in mAP on the ICFG-PEDES dataset, as observed in No. 0 vs. No. 1. Additionally, all evaluation metrics on CUHK-PEDES and ICFG-PEDES degrade, further confirming the superiority of A-SDM. Moreover, when combined with the EFA module, the advantage of A-SDM becomes even more pronounced. As shown in No. 2 vs. No. 3, replacing the A-SDM module with the SDM module results in 0.43% and 0.48% decrease in Rank-1 and Rank-5 on the CUHK-PEDES dataset, respectively, as well as a 1.55% drop in Rank-5 and a 0.58% decline in mAP on the RSTPReid dataset. These results collectively validate the consistent and significant impact of A-SDM to performance.
Table 6. Ablation analysis of FMFA modules across three public benchmarks.
@@ 4.3. Ablation Study
SDM A-SDM EFA Rank-1 Rank-5 Rank-10 mAP Rank-1 Rank-5 Rank-10 mAP Rank-1 Rank-5 Rank-10 mAP
0 Baseline $\checkmark$ 73.45 89.38 93.69 66.13 63.48 80.16 85.78 38.20 59.50 81.80 88.85 47.44
1 +A-SDM $\checkmark$ 74.04 89.86 93.89 66.45 64.26 80.59 85.90 39.39 60.25 81.45 88.70 47.69
2 +EFA $\checkmark$ $\checkmark$ 73.73 89.64 94.04 66.40 63.77 80.39 85.86 39.17 60.45 82.30 89.25 47.64
3 FMFA $\checkmark$ $\checkmark$ 74.16 90.12 94.10 66.66 64.29 80.48 85.93 39.43 61.05 83.85 89.80 48.22
Effect of The EFA Module. To improve the model’s global matching performance, the Explicit Fine-grained Alignment (EFA) module introduces fine-grained cross-modal interaction based on a sparse similarity matrix. The impact of the EFA module is illustrated by comparing the results of No. 0 vs. No. 2 and No. 1 vs. No. 3. Specifically, as shown in No. 1 vs. No. 3, removing the EFA module from FMFA leads to a performance drop of 0.26% and 0.21% in Rank-5 and Rank-10 on the CUHK-PEDES dataset, and a more significant decline of 2.40% and 1.10% in Rank-5 and Rank-10, along with a 0.53% decrease in mAP on the RSTPReid dataset. However, EFA causes a 0.11% drop in Rank-5 on ICFG-PEDES, suggesting that its sparse and hard coding strategy, which focuses only on the most relevant patches, may overlook other informative ones and lead to information loss. Notably, this comparison also reflects the joint ablation of EFA and the A-SDM module, further verifying the complementary effect between the two modules. Moreover, to further validate the individual contribution of the EFA module, we design an additional experiment that retains the SDM module while removing only the EFA module, as observed in No. 0 vs. No. 2. In this setting, the absence of EFA results in 0.38%, 0.29%, and 0.95% drops in Rank-1 on CUHK-PEDES, ICFG-PEDES and RSTPReid, and causes a 1.23% drop in mAP on ICFG-PEDES. These evaluations highlight the effectiveness of the EFA module.
4.4. Parameter Study
We perform a parameter study on the CUHK-PEDES dataset, examining three hyperparameters—the weight factor $\alpha$, the sparsity threshold $\sigma$, and the factor $\lambda$—as well as the contribution weights of the proposed loss functions, $\mathcal{L}_{A\text{-}SDM}$ and $\mathcal{L}_{EFA}$. When examining a specific parameter, all other parameters are maintained as specified in Section 4.1.
@@ 4.4. Parameter Study
Hyperparameters Analysis. As shown in Figure 4(a), we vary the weight factor $\alpha$ from 1 to 20. The results on CUHK-PEDES show that setting $\alpha$ to 10 achieves the highest performance across the evaluated metrics, suggesting that the adaptive pull force on unmatched positive pairs is optimal. However, when $\alpha$ is increased to 20, the performance drops significantly because the pull force becomes excessively strong, causing unmatched positive pairs to fail to align properly and instead over-pulling mismatched positives, which leads to false positives. We further vary the sparsity threshold $\sigma$ from $0.1/N$ to $20/N$ and the LSE factor $\lambda$ from 0.1 to 10, where $N$ denotes the number of image patches. As illustrated in Figure 4(b) and Figure 4(c), setting $\sigma$ to $1/N$ and $\lambda$ to 1 yields the best overall performance. When $\sigma$ is set too high, only a few highly relevant patches are retained, leading to the loss of semantic information, whereas an excessively low $\sigma$ preserves most patches and weakens the ability to capture discriminative features. According to Eq. (10), a large $\lambda$ emphasizes the differences among patch responses and makes the pooling operation less robust to noise, while a small $\lambda$ smooths these responses excessively and reduces feature discrimination. Moreover, to evaluate the stability of these hyperparameters, we further set $\sigma$ around $1/N$ (i.e., $0.9/N$ and $1.1/N$) and $\lambda$ around 1 (i.e., 0.9 and 1.1). The results show that although slightly adjusting $\sigma$ and $\lambda$ leads to a minor performance drop, our model still obtains high performance, demonstrating the robustness of our method.
Table 7. The sensitivity analysis of the weight of $\mathcal{L}_{A-sdm}$ and $\mathcal{L}_{efa}$ on the CUHK-PEDES dataset, where one weight is varied and the other is fixed at 1.0.
@@
Table 8. Ablation study verifying the necessity of the constant term “+1” in A-SDM on the CUHK-PEDES dataset.
Setting “+1” Term Rank-1 Rank-5 Rank-10
A-SDM w/o “+1” 26.12 48.36 59.18
A-SDM w “+1” $\checkmark$ 74.16 90.12 94.10
Loss Function Weights. We perform experiments using the CUHK-PEDES dataset to investigate the influence of the weights of the two proposed loss functions, $\mathcal{L}_{A-sdm}$ and $\mathcal{L}_{efa}$, varying them from 0.1 to 10.0, as shown in Table 7. To comprehensively explore the optimal combination of the proposed loss weights, we only vary one weight at a time while keeping the other fixed at 1 to ensure a controlled comparison. The results show that setting both weights to 1 yields the best performance. Notably, increasing the weight of $\mathcal{L}_{efa}$ to 5.0 or 10.0 results in gradient explosion, making the model untrainable and causing all metrics to drop below 1.
@@
The “+1” term in A-SDM. We carry out experiments on the CUHK-PEDES dataset to validate the necessity of the constant “+1” in A-SDM. As defined in Eq. (2), the constant term “+1” ensures a positive lower bound for the adaptive weight, preventing it from becoming zero when a pair is correctly matched. Without this term, gradients of correctly matched positive pairs diminish, leading to unstable optimization and degraded performance during training. Table 8 indicates that omitting the “+1” leads to a notable decline in retrieval accuracy, illustrating that the effectiveness of model is reduced. Therefore, the “+1” term is a crucial component that ensures stable optimization and consistent performance of A-SDM.
@@ 4.5. Qualitative Results
Visualization of The Sparse Process. We visualize the similarity maps before and after the sparse process (i.e., Eq. (5) and Eq. (6)) in the EFA module, as shown in Figure 5. Patches with high similarity (greater than 0.75) are preserved, as indicated by the black frames in the left part of Figure 5. In contrast, patches with low similarity (below 0.5) are suppressed to values under 0.25 and thus omitted during aggregation, as shown by the pink frames on the left. When most patches in an image exhibit relatively high similarity (around 0.75), the sparse process retains only those with the highest similarity while reducing the similarity of the remaining patches below 0.5, highlighted by the purple frames in the right part of Figure 5. This guarantees that the subsequent aggregation emphasizes only the most pertinent patches, decreasing computational and memory overhead while preserving performance. Unlike implicit aggregation methods based on attention mechanisms, EFA explicitly aggregates image patches and text tokens, allowing us to observe whether the most relevant patches are effectively aggregated by visualizing the similarity between patches and tokens.
Inference Time Comparison. We compare the inference time between FMFA and recent matching methods (e.g., PLOT (park2024plot) and RaSa (bai2023rasa)) on the test sets of three datasets, as shown in Table 9. As a global matching method, FMFA only computes global features during inference, thus achieving a higher inference speed than local matching methods. Even when compared with the recent global matching method DM-Adapter (liu2025dm), FMFA still achieves consistently faster inference across all three datasets. Moreover, the efficiency advantage of FMFA becomes more pronounced with the growth of the test set. For instance, the inference time of FMFA vs. PLOT increases from 3s vs. 5s on RSTPReid to 50s vs. 91s on ICFG-PEDES. These comparisons clearly demonstrate that FMFA achieves superior inference speed compared with recent methods.
Table 9. Comparison of the inference time (s) between FMFA and recent methods.
@@
DM-Adapter (liu2025dm) 8 16 78
PLOT (park2024plot) 5 16 91
APTM (yang2023towards) 10 29 95
SCVD (wei2024fine) - 191 1369
RaSa (bai2023rasa) 388 1168 3871
@@
Figure 6. Top-5 retrieval results for each text query on CUHK-PEDES, comparing baseline with FMFA. Target image, correct matches, and mismatches are outlined in black, green, and red.
Visualization of Top-5 Retrieval Results. Figure 6 compares the top-5 retrieval results between the baseline IRRAR and our proposed FMFA on the CUHK-PEDES dataset. Figure 6 illustrates that FMFA delivers more precise retrieval results, correctly identifying images that the baseline fails to match. For query texts where the baseline performs well, FMFA further improves performance by retrieving more relevant pedestrian images (e.g., Figure 6(a) and Figure 6(c)). Even for hard negative samples, where the baseline struggles to retrieve the correct image, FMFA still enhances the similarity between positive pairs (e.g., Figure 6(b) and Figure 6(d)). This is because our proposed FMFA focuses on the unmatched positive pairs and adaptively pulls the positive pairs closer. More comparisons of the top-5 retrieved results are provided in Appendix B.
5. Conclusion
In summary, we propose a cross-modal Full-Mode Fine-grained Alignment (FMFA) framework to learn discriminative global text-image representations through full-mode fine-grained alignment, including explicit fine-grained alignment and existing implicit relational reasoning. We design an Adaptive Similarity Distribution Matching (A-SDM) module to concentrate on unmatched positive pairs, adaptively pulling them closer. In addition, to achieve cross-modal fine-grained alignment, we introduce an Explicit Fine-grained Alignment (EFA) module, which explicitly aggregates local text and image representations based on the sparse similarity matrix and employs a hard coding method. These modules function together to project images and text into a shared embedding space. Comprehensive experiments across three datasets confirm the effectiveness and superior performance of our FMFA framework.
Limitation. The fixed threshold in the sparse process only keeps the most relevant patches, which may result in the loss of semantic information and limit the effective aggregation of local features, thereby affecting the overall performance of our model. Incorporating adaptive methods that capture complete semantic information (e.g., tree transformer (wang2019tree)) could further enhance our model.
Acknowledgements.
This work was supported by the National Natural Science Foundation of China (No. 62302080), Guangxi Key Research and Development Program (No. Guike AB24010112), National Foreign Expert Project of China (No. S20240327), Sichuan Science and Technology Program (No. 2025HJRC0021) and Sichuan Province Innovative Talent Funding Project for Postdoctoral Fellows (No. BX202312).
Appendix A Theoretical Analysis of Hard Coding in EFA
Given visual embeddings of an image $\boldsymbol{V}\in\mathbb{R}^{N\times d}$ and its joint embeddings $\boldsymbol{E}\in\mathbb{R}^{L\times d}$, where $N$ is the number of image patches and $L$ is the length of the joint embeddings, both the hard coding method and the soft coding method calculate the similarity matrix $A$ to obtain the final similarity between $\boldsymbol{V}$ and $\boldsymbol{E}$, which requires the same time complexity $\mathcal{O}(NLd)$. When processing the similarity matrix $A$, our hard coding method calculates the maximum similarity for each row, which has a time complexity of $\mathcal{O}(N)$, as described in Eq. (9) and Eq. (10). However, the soft coding method processes all similarities, resulting in a time complexity of $\mathcal{O}(NL)$. Moreover, the spatial complexity of the hard coding method is $\mathcal{O}(NL)$, which is significantly lower than the spatial complexity of the soft coding method, $\mathcal{O}(Ld)$. This is based on the fact that $N\ll d$. Therefore, our hard coding method can effectively reduce the cost of training time and memory.
Appendix B More Comparisons of The Top-5 Retrieved Results
Figure 7 and Figure 8 show the comparisons of the top-5 retrieved results between the baseline and FMFA on the ICFG-PEDES and RSTPReid datasets, respectively. These comparisons further highlight the superiority and effectiveness of the proposed FMFA framework.
@@ Appendix B More Comparisons of The Top-5 Retrieved Results
Figure 7. Top-5 retrieval results for each text query on ICFG-PEDES, comparing baseline with FMFA. Target image, correct matches, and mismatches are outlined in black, green, and red.
@@
Figure 8. Top-5 retrieval results for each text query on RSTPReid, comparing baseline with FMFA. Target image, correct matches, and mismatches are outlined in black, green, and red.